OpenStudy (anonymous):

fid the y- intercept of 6x-y=-12

OpenStudy (angela210793):

-y=-12-6x y=6x+12 y-intercept 12

OpenStudy (anonymous):

6x-y=-12 6(0)-y=-12 -y=-12 -1 -1 y=12 The y-intercept is 12.